The monthly incomes for 12

randomly selected​ people, each with a​ bachelor's degree in​ economics, are shown on the right. Complete parts​ (a) through​ (c) below.
Assume the population is normally distributed.

4450.56
4596.08
4366.27

4455.71
4151.02
3727.27
4283.79
4527.36
4407.24
3946.73
4023.58
4221.79
Question content area bottom
Part 1
​(a) Find the sample mean.
x overbar
equalsenter your response here
​(Round to one decimal place as​ needed.)

Answer by math_tutor2020(3817)


To find the sample mean, add up the values and then divide by the sample size.

Example:
The sample mean of {1,2,7} is (1+2+7)/3 = 10/3 = 3.33 approximately

I recommend pasting the given data into a spreadsheet. Then you can add the values quickly with the command named "SUM".
Another command is called "AVERAGE" to directly compute the sample mean.

For example, if your data values are in cells A1 through A12, then:
=SUM(A1:A12) will compute the sum of the items in cells A1 through A12
=AVERAGE(A1:A12) will compute the arithmetic mean of the items in cells A1 through A12
Do not forget about the equal sign up front. It is needed to evaluate the command. Otherwise, the spreadsheet will interpret it as plaintext.

I'll let the student do the computations I have described.